OpenStudy (anonymous):

Let the function, f, be defined by f(1)=2 and f(n)=n^2+f(n-1) for all natural numbers n (greater or equal to) 2. Find the value of f(5)

OpenStudy (anonymous):

f(1)=2 then, f(2)=2^2+ f(1) f(3)=3^2+ f(2) f(4)=4^2+ f(3) f(5)=5^2+ f(4) need to evaluate f(n) for every n<5 first before u can calculate f(5).